OpenStudy (anonymous):

Solve. Clear fractions first 2/5x-1/3x=4

OpenStudy (anonymous):

when i cleared the fractions i got 1/15x = 4

OpenStudy (anonymous):

2/5x - 1/3x = 4 Put the fractions over a common denominator, 15x (6-5)/15x = 4 1 = 60x x = 1/60

OpenStudy (anonymous):

Oh since you already got to 1/15x = 4, just multiply the 15x over to get 1 = 60x, and then divide both sides by 60

OpenStudy (anonymous):

oh k now what if the problem is 3/4 + 1/5x = 1

OpenStudy (anonymous):

when i cleared the fractions i got 19/20x